# HiL-Bench (Human-in-Loop Benchmark) Do Agents Know When to Ask for Help?

### 10 Apr 2026 v1

@@ 1 Introduction
Not only do current benchmarks actively select against detecting this failure mode, but they also actively encourage against detecting it. SWE-Bench (and its derivations) (Jimenez et al. (2024), Chowdhury et al. (2024), Deng et al. (2025)), HumanEval (Chen et al., 2021), and BIRD-SQL Li et al. (2024) supply fully specified tasks and reward confident autonomy: an agent that silently assumes its way past a missing requirement and get lucky will score identically to one that would have asked for clarification. The result is a misleading feedback loop: high benchmark scores encourage deployment, deployment surfaces the judgment failures benchmarks never tested, and the >>90% pilot failure rate persists because the evaluation and the failure mode occupy different worlds (Egg et al., 2025), presenting a façade of readiness.
@@
    Ask-F1, a precision-recall metric that distinguishes the ability to detect context gaps from the judgment of when to ask.
@@ 2 Related Work
The dominant paradigm of benchmarks supplies complete specifications and scores silent execution. Function-level benchmarks (Chen et al., 2021; Austin et al., 2021; Jain et al., 2024), repository-scale code repair (Jimenez et al., 2024; Deng et al., 2025), and longer-horizon agent evaluations spanning workplace tasks, web navigation, GUI interaction, and tool usage (Xu et al., 2024; Zhou et al., 2023; Drouin et al., 2024; Xie et al., 2024; Bandi et al., 2026) all share one property: they make it impossible to distinguish an agent that guesses its way past an unclear requirement from one that asks, because their environment and task composition do not present opportunities where such agents would showcase their different behaviors. No existing benchmark penalizes confident wrongness born from incomplete information, only wrongness in general.
@@
Prior work tests whether models can identify ambiguity across dialogue (Aliannejadi et al., 2020; Rao & III, 2018; Majumder et al., 2021), question answering (Min et al., 2020; Xu et al., 2019), constraint satisfaction (Li et al., 2025), multi-turn simulation (Qian et al., 2025), and varied ambiguity types (Suri et al., 2025; Zhang et al., 2024). Recent efforts approach agentic settings: Wang et al. (2025) evaluate tool-use agents under unclear instructions; Andukuri et al. (2024) train LMs to generate clarifying questions; and Huang et al. (2025b), the closest prior work, constructs tasks with multiple implicit blockers and applies RL to improve downstream performance. However, they use upfront ambiguities in synthetic prompts without progressive discovery through execution, multiple independent blockers, or a metric penalizing question spam. A parallel thread converts static coding tasks into multi-turn collaborative processes (Pan et al., 2025; Zhang & Choi, 2025; Lahiri et al., 2022; Nijkamp et al., 2022) or delegates clarification to peer agents (Huang et al., 2023; Qian et al., 2024).
@@ 3.2 Blocker Design
The three blocker types correspond to distinct classes of production failures. Missing information (42% of blockers are required values absent from the specification: a parser must return a fallback on failure but the spec never names it; a SQL question requires counting “quick pit stops” but no duration threshold is defined. Ambiguous requests (36%) admit multiple valid implementations or interpretations: epoch segments in version strings must be handled “appropriately” but multiple consistent strategies exist (strip, normalize, delegate); filtering for “countries in the Middle East” can have different inclusions depending on the user. Contradictory information (22%) present specifications that cannot both be satisfied: one requirement says admin roles have a special privilege, while guidance from another source says otherwise; a SQL question requests statistics about northern California schools but specifically names southern California ones. The full taxonomy with additional examples appears in Appendix A.
@@ 4.2 Results: The Judgment Gap
The central finding is a large per-domain judgment gap (Tab. 1). Under full information, models reach 86–91% pass@3 on SQL and 64–88% on SWE, but once they must decide when to invoke ask_human(), the best model reaches only 38% on SQL and 12% on SWE. Ask-F1 averages 40.5% on SQL and 37.4% on SWE. Because no task can be passed without resolving every blocker, this collapse confirms that judgment, not raw capability, is the binding constraint.

### 13 Apr 2026 v2

@@ 1 Introduction
Not only do current benchmarks actively select against detecting this failure mode, but they also actively encourage against detecting it. SWE-Bench (and its derivations) (Jimenez et al. (2024), Chowdhury and others (2024), Deng et al. (2025)), HumanEval (Chen et al., 2021), and BIRD-SQL Li et al. (2024) supply fully specified tasks and reward confident autonomy: an agent that silently assumes its way past a missing requirement and get lucky will score identically to one that would have asked for clarification. The result is a misleading feedback loop: high benchmark scores encourage deployment, deployment surfaces the judgment failures benchmarks never tested, and the >>90% pilot failure rate persists because the evaluation and the failure mode occupy different worlds (Egg et al., 2025), presenting a façade of readiness.
@@ 2 Related Work
The dominant paradigm of benchmarks supplies complete specifications and scores silent execution. Function-level benchmarks (Chen et al., 2021; Austin et al., 2021; Jain et al., 2024), repository-scale code repair (Jimenez et al., 2024; Deng et al., 2025), and longer-horizon agent evaluations spanning workplace tasks, web navigation, GUI interaction, and tool usage (Xu et al., 2024; Zhou and others, 2023; Drouin et al., 2024; Xie et al., 2024; Bandi et al., 2026) all share one property: they make it impossible to distinguish an agent that guesses its way past an unclear requirement from one that asks, because their environment and task composition do not present opportunities where such agents would showcase their different behaviors. No existing benchmark penalizes confident wrongness born from incomplete information, only wrongness in general.
@@
Prior work tests whether models can identify ambiguity across dialogue (Aliannejadi et al., 2020; Rao and III, 2018; Majumder et al., 2021), question answering (Min et al., 2020; Xu et al., 2019), constraint satisfaction (Li et al., 2025), multi-turn simulation (Qian et al., 2025), and varied ambiguity types (Suri and others, 2025; Zhang et al., 2024). Recent efforts approach agentic settings: Wang et al. (2025) evaluate tool-use agents under unclear instructions; Andukuri et al. (2024) train LMs to generate clarifying questions; and Huang et al. (2025b), the closest prior work, constructs tasks with multiple implicit blockers and applies RL to improve downstream performance. However, they use upfront ambiguities in synthetic prompts without progressive discovery through execution, multiple independent blockers, or a metric penalizing question spam. A parallel thread converts static coding tasks into multi-turn collaborative processes (Pan and others, 2025; Zhang and Choi, 2025; Lahiri et al., 2022; Nijkamp et al., 2022) or delegates clarification to peer agents (Huang et al., 2023; Qian et al., 2024).
@@ 3.2 Blocker Design
The three blocker types correspond to distinct classes of production failures. Missing information (42%) of blockers are required values absent from the specification that surface only during execution: e.g., an unspecified parser fallback value on failure; a SQL question requires counting “quick pit stops” but no duration threshold is defined. Ambiguous requests (36%) admit multiple valid implementations or interpretations: epoch segments in version strings must be handled “appropriately” but multiple consistent strategies exist (strip, normalize, delegate); filtering for “countries in the Middle East” can have different inclusions depending on the user. Contradictory information (22%) present specifications that cannot both be satisfied: one requirement says admin roles have a special privilege, while guidance from another source says otherwise; a SQL question requests statistics about northern California schools but specifically names southern California ones. The full taxonomy with additional examples appears in Appendix A.
@@ 4.2 Results: The Judgment Gap
The central finding is a large per-domain judgment gap (Table. 1). Under full information, models reach 86–91% pass@3 on SQL and 64–88% on SWE, but once they must decide when to invoke ask_human(), the best model reaches only 38% on SQL and 12% on SWE. Ask-F1 averages 40.5% on SQL and 37.4% on SWE. Because no task can be passed without resolving every blocker, this collapse confirms that judgment, not raw capability, is the binding constraint.